Telomere Length Estimation

Objective

To estimate the length of telomeres on your extracted gDNA. Background Telomeres are repetitive nucleotide elements at the ends of chromosomes that protect chromosomes from degradation and genetic information loss. Normal diploid cells lose telomeres with each cell cycle. Telomere length, therefore, decreases over time and may predict lifespan. Telomere shortening has negative effects on health conditions and has been linked to many health issues including aging and cancer. Accurate and consistent quantification of telomere length is important in many aspects of cell biology such as chromosomal instability, DNA repair, senescence, apoptosis, cell dysfunctions, and oncogenesis.

Just like SNPs, telomere length is an aspect of an individual’s genome that is highly unique and personal, and therefore this characteristic has also been exploited by biotechnology companies such as teloyears.com, or lifelength.com. These companies offer personalized genomic testing aimed at estimating telomere length and comparing each individual result to a growing database. Scientific evidence is growing showing that the telomere shortening process may be reversible, or at least, preventable, therefore contributing to increase a person’s life expectancy. As indicated above, abnormally short telomeres may also be a biomarker of disease.

From a technical standpoint, telomere length estimation relies on qPCR, by using primers designed to anneal to the highly-repeated motifs that make up telomeres. The reaction also includes a primer for a reference gene (Single Copy Reference, or SCR) that recognizes and amplifies a 100 bp-long region on human chromosome 17, and serves as reference for data normalization. Longer telomeres mean higher concentrations of template for the qPCR reactions, and results in a higher concentration of PCR products (obtained with fewer cycle) when compared to the one-copy reference.

Overview of the method

We will use a classic SYBR green based PCR set-up a qPCR reaction. Your gDNA will be used as a template for two separate reactions that include either the telomere primers or the SCR primers. This time, we will NOT add a non-template control (NTC)but we will include a POSITIVE CONTROL that contains a known amount of telomeres (estimated at 233kb). Please note that, this time, your gDNA is included as part of the Master Mix. The primers are the variable in these particular reactions.

Post-Lab Calculations


1. For telomere (TEL), ∆Cq (TEL) is the quantification cycle number difference of TEL between the target and the reference genomic DNA samples.

∆Cq (TEL) = Cq (TEL, target sample) - Cq (TEL, reference sample)

Note: the value of ∆Cq (TEL) can be positive, 0, or negative.

2. For single copy reference (SCR), ∆Cq (SCR) is the quantification cycle number difference of SCR between the target and the reference genomic DNA samples.


∆Cq (SCR) = Cq (SCR, target sample) - Cq (SCR, reference sample)


Note: the value of ∆Cq (SCR) can be positive, 0, or negative.


3. ∆∆Cq = ∆Cq (TEL) - ∆Cq (SCR)

4. Relative telomere length of the target sample to the reference sample (fold) = 2-∆∆Cq

5. The total telomere length of the target sample. = Reference sample telomere length x 2-∆∆Cq

HOW TO GET Cq?

The quantitative cycle (Cq), also referred to as the threshold cycle (Ct) is the number of cycles required for the fluorescent to be detected and exceed the background level. It corresponds to the number of cycles at which the exponential curve associated with qPCR starts. These numbers are obtained once the reactions are performed and are part of the output provided by the qPCR machine.
Refer to the available file on Canvas to retrieve the Cq values associated with our reactions. Notice that the program automatically recognizes replicate reactions and calculate the average Cq values for our reference reactions.
Locate your own Cq values, and perform the calculations as indicated above.

The reference sample telomere length is 233kb per diploid cell, so you should be able to calculate your own telomere length per diploid cell

LAB RESULTS:
· REFERENCE
o SCR 23.45
o TEL 14.78
· gDNA (my genome results/target sample)
o SCR 22.05
o TEL 17.22


PLEASE HELP TO ANSWER THESE QUESTIONS AND SHOW STEPS

1. Based on your understanding of qPCR and Cq values, explain why the SCR Cq values are higher than the TEL values (on average)

2. Perform the calculations of gDNA Cq values, and indicate the estimated telomere length per diploid cells

3. How many telomeres (chromosome ends) gDNA have in each of your diploid cells? Use this number to calculate the average length of each of your individual telomeres (chromosome ends).

4. In humans the telomere motif is TTAGGG. How many motifs do each of gDNA chromosome ends contain (on average)?
Answer #1

1) SCR is a 100 bp long region in human chromosome no 17. It acts as a reference, that means it acts as control. It has only one copy number in each cell. The SCR Cq values are higher than the TEL values, as the copy number of your gene of interest (TEL) is more than one.

2) ∆Cq (TEL) = Cq (TEL, target sample) - Cq (TEL, reference sample)

So, ∆Cq (TEL) = 17.22 - 14.78 = 2.44

∆Cq (SCR) = Cq (SCR, target sample) - Cq (SCR, reference sample)

So, ∆Cq (SCR) = 22.05 - 23.45 = -1.4

∆∆Cq = ∆Cq (TEL) - ∆Cq (SCR)

So, ∆∆Cq = 2.44 - (-1.4) = 2.44 + 1.4 = 3.84

Relative telomere length of the target sample to the reference sample (fold) = 2-∆∆Cq

So, Relative telomere length of the target sample to the reference sample (fold) = 2^-3.84 = 0.07

The total telomere length of the target sample. = Reference sample telomere length x 2-∆∆Cq

So, the total telomere length of the target sample. = 233 X 0.07 = 16.31 kb

3) If this diploid cell is extracted from a human, there are 46 telomeres in each of your diploid cell. So, as the total telomere length is 16.31 kb, the average length of each of your individual telomere is (16.31/46) = 0.35 Kb = 350 bp.

4) In humans the telomere motif is TTAGGG. So, each motif contains 6 nucleotides. So, each of gDNA chromosome ends contain (350/6) = 58 motifs.